### v2

# Schedule On the Fly: Diffusion Time Prediction for Faster and Better Image Generation

Abstract
Diffusion and flow models have achieved remarkable successes in various applications such as text-to-image generation. However, these models typically rely on the same predetermined denoising schedules during inference for each prompt, which potentially limits the inference efficiency as well as the flexibility when handling different prompts. In this paper, we argue that the optimal noise schedule should adapt to each inference instance, and introduce the Time Prediction Diffusion Model (TPDM) to accomplish this. TPDM employs a plug-and-play Time Prediction Module (TPM) that predicts the next noise level based on current latent features at each denoising step. We train the TPM using reinforcement learning, aiming to maximize a reward that discounts the final image quality by the number of denoising steps. With such an adaptive scheduler, TPDM not only generates high-quality images that are aligned closely with human preferences but also adjusts the number of denoising steps and time on the fly, enhancing both performance and efficiency. We train TPDMs on multiple diffusion model benchmarks. With Stable Diffusion 3 Medium architecture, TPDM achieves an aesthetic score of 5.44 and a human preference score (HPS) of 29.59, while using around 50% fewer denoising steps to achieve better performance. We will release our best model alongside this paper.
footnotetext: The project was initiated and supported by MAPLE Lab at Westlake University.footnotetext: $\dagger$ The work was done during an internship at MAPLE lab.footnotetext: Corresponding author.
1 Introduction
In recent years, deep generative models, including diffusion models [52, 13, 50] have achieved extraordinary performance across a variety of tasks, including image synthesis [47, 44, 43, 18, 19], video generation [14, 4, 63], and others [22, 40, 49, 36]. As a multi-step denoising framework, diffusion models progressively refine random noise into coherent data through iterative sampling, which underlies their impressive capabilities in generating high-quality, diverse outputs.
However, inference with a diffusion model always involves manually selecting a noise scheduler, e.g. how the noise level changes step by step when denoising a clean image from Gaussian noise. This requires the user to delicately adjust parameters like the number of steps, resulting in a high threshold for customers to use. Numerous works attempt to find the optimal scheduler to liberate the users from these tedious tasks.
Figure 1: Samples generated by TPDM-FLUX.1-dev showcase stunning visual effects while adaptively adjusting inference steps based on the target output. The number in the lower right corner of each image indicates the inference steps used.
The leading flow-based models, Stable Diffusion 3 and Flux, provide recommended schedulers that adjust noise levels only based on the targets’ resolution. Sabour [46] and Xia [60] explore ways to fine-tune the schedule for a model, improving either its efficiency or overall performance. In addition, some new one-step generators [16, 33, 34, 35] also achieve impressive performance. Despite the excellent performance they achieve, most of these works hold the assumption that there exists a universally applicable schedule that is optimal for all prompts and images, which is doubtful.
Take several images in Fig. 4 as an example. Images on the right are rich in content, requiring more denoising steps to capture finer details. In contrast, the images on the left are relatively simple and can be generated with fewer steps without compromising quality. Moreover, Karras [20] also proves that different noise schedules also greatly affect the generation quality. Thus, we may ask: Is it possible to adaptively adjust both the number of denoising steps and the noise level at each step during inference, without any manual intervention from the user?
In this paper, we propose Time Prediction Diffusion Models(TPDMs) that can adaptively adjust both the number of steps and the denoising strength during inference. This is achieved by implementing a plug-and-play Time-Prediction Module (TPM) that can predict the next diffusion time conditioned on the latent features at the current step, so that the noise schedule can be adjusted on the fly.
TPM is trained with reinforcement learning. We regard the multi-step denoising process as an entire trajectory, and the quality of the generated image discounted by the number of steps as the reward. The image quality is measured by a reward model aligned with human preference [62].
TPM can be easily integrated into any diffusion models with marginal additional computation, and allow them to automatically adjust hyper-parameters like sample steps and the noise level at each step, reaching a balance between image quality and efficiency without human interventions. Moreover, during training, the model rolls out the diffusion process the same as in inference, directly optimizing the inference performance and reducing denoising steps.
We implement TPDM on several start-of-the-art models, including Stable Diffusion and Flux. With an adaptive noise schedule, our model can also generate $50\%$ fewer steps on average with image quality on par or slightly better (0.322 CLIP-T, 5.445 Aesthetic Score, 22.33 Pick Score, 29.59 HPSv2.1) than Stable Diffusion 3. These results demonstrate that TPDM has the potential to either pursue high-quality image generation or improve model efficiency.
Our contributions are summarized below:
  • We introduce the Time Prediction Diffusion Model (TPDM) that can adaptively adjust the noise schedule during inference, achieving a balance between image quality and model efficiency.
  • To train TPDM, we maximize the generated image quality discounted by the number of steps with reinforcement learning, directly optimizing final performance and efficiency.
  • Our model demonstrates improved performance across multiple evaluation benchmarks, achieving better results with fewer inference steps.
2 Related Works
2.1 Diffusion Models
Diffusion Probability Models (DPMs) [13, 50] recover the original data from pure Gaussian noise by learning the data distribution across varying noise levels. With their strong adaptability to complex data distributions, diffusion models have achieved remarkable performance in diverse fields such as images [39, 40, 42, 15, 21], videos [14, 4, 63], and others [65, 57, 9, 64], significantly advancing the capabilities of Artificial Intelligence Generated Content.
2.2 Noise Schedule
In order to generate an image, the model must determine the diffusion time for each step. This can be achieved using either discrete-time schedulers [13, 51] or continuous-time schedulers [29, 26] depending on the model. Typically, the diffusion time reflects the noise strength at each step, and most existing approaches rely on pre-determined schedules. Currently, the leading flow-based models, Stable Diffusion 3[10] and FLUX[23], provide recommended schedulers that adjust noise levels only based on the target resolution.
There exist some methods fine-tuning the scheduler to speed up sampling or improve image quality. Xia et al. [60] predict a new denoising time for each step to find a more accurate integration direction, Sabour et al. [46] uses stochastic calculus to find the optimal sampling plan for different schedulers and different models. Wang et al. [56] leverages reinforcement learning to automatically search for an optimal sampling scheduler. Also, adjust the schedule is also effective is also effective in other areas such as molecular generation [17]. Some one-step generators[34, 35, 16] with diffusion distillation [31] also achieve impressive performance.
All of the above noise schedulers not only have high thresholds that require users to adjust many parameters, but also use the same denoising schedule for all prompts and images. On the contrary, TPDM can adaptively adjust the noise schedule during inference, and dynamically select the optimal sampling plan with decreased number of sampling steps for all prompts, achieving a balance between image quality and model efficiency. We will introduce the detailed definitions and practical algorithms of TPDM in Sec. 3.
2.3 Reinforcement Learning and Learning from Human Feedback
Reinforcement Learning from Human Feedback (RLHF) has recently gained significant attention in the field of large language models (LLMs) [48, 1, 53] and is gradually expanding into other domains. Advances in diffusion models have increasingly incorporated reward models to enhance alignment with human preferences [25, 62, 59]. By treating latents as actions, diffusion models can be optimized with policy gradient (DDPO [3], DPOK[12]), actor-critic framework (DACER [58]), direct preference optimization (Diffusion-DPO [55]) and other reinforcement algorithms. Some recent works have also studied the RLHF for few-step generative models [32, 37]. Since these approaches rely on the Gaussian reverse processes of the diffusion SDE sampler, however, they are challenging to apply in flow models. On the contrary, in this paper we treat the diffusion time as the action instead of latents. Our goal is to adjust the schedule to achieve better quality with fewer steps, offering a more general and flexible solution.
3 The Proposed Approach
In this section, we first provide a brief review of the fundamental principles of diffusion models, followed by an introduction to the Time Prediction Module (TPM). Finally, we detail the training algorithm for the TPM.
Figure 2: The Inference Process of TPDMs: The horizontal axis represents diffusion time, ranging from 1 to 0. The image starts from random noise $x_{t_{0}}$ and is progressively denoised until a clean image $x_{t_{N}}$. Meanwhile, the reward is calculated for the final image and discounted by $\gamma$ to influence previous steps.
Figure 3: The architecture of TPDM involves a frozen Diffusion Models, and a plug-and-play Time Prediction Module.
3.1 Preliminary
Diffusion Models learn to generate images with a reverse process that gradually removes noise from a sample. The leading paradigm for implementing this reverse process is flow matching [27, 10, 11]. Thus we introduce how flow-matching models work, and the detailed structure inside current start-of-the-art models here.
We consider a generative model that establishes a mapping between samples $x_{1}$ drawn from a noise distribution $p_{1}$ and samples $x_{0}$ from a data distribution $p_{0}$. The Flow Matching objective aims to directly regress a vector field $v_{t}$, which generates a probability flow, enabling the transition from $p_{1}$ to $p_{0}$.
$\mathcal{L}_{FM}(\theta)=\mathbb{E}_{t,p_{t}(x)}\left\|v_{\theta}(x_{t},t)-u(x_{t},t)\right\|^{2}$ (1)
The flow-matching model with parameters $\theta$ aims to predict the noise prediction function $v_{\theta}(x_{t},t)$, which approximates the true velocity field $u(x_{t},t)$ that guides the diffusion process from the noise distribution to the clean distribution of generated samples. Thus, we can get the diffusion ODE:
$\frac{dx_{t}}{dt}=v_{\theta}(x_{t},t)$ (2)
During inference, suppose we generate an image with $N$ steps, each step has a time $t_{n}$ corresponding to its noise strength. Then, the $n$-th generation step can be formulated as:
$x_{t_{n+1}}=x_{t_{n}}+(t_{n+1}-t_{n})\cdot v_{\theta}(x_{t_{n}},t_{n})$ (3)
Usually, in a typical flow matching algorithm, there is a pre-determined schedule that arranges $t_{n}$ every step.
Currently, many state-of-the-art diffusion models are built upon the DiT architecture [24, 8, 10], which only focus on conditional image generation and employ a modulation mechanism to condition the network on both the diffusion time during the diffusion denoising step and the text prompts.
With such a model, we can perform a single denoising step given the latent feature and current noise level.
3.2 Time Prediction Diffusion Model (TPDM)
As aforementioned, we need a series of denoising steps to generate an image with the trained diffusion model. Usually, a fixed noise schedule is applied in this process for all prompts, assigning a pre-determined noise level for each step.
On the contrary, to enable the model to adjust the noise schedule on the fly, the TPDM predicts how noise strength $t$ decays between adjacent steps. This ensures the noise strength is monotonously decreasing to reflect the denoising progress, avoiding backward progression. To be compatible with training in Sec. 3.3, TPM estimates the distribution over the denoising schedule, rather than predicting an exact value. Suppose we are performing the $n$-th denoising step. Besides the denoising outputs at the current step, TPDM predicts the distribution of the decay rate $r_{n}$ as well. This distribution can be parameterized as a Beta Distribution over $(0,1)$, where the TPDM needs to estimate two parameters $\alpha$ and $\beta$. We notice that when $\alpha>1$ and $\beta>1$ it can ensure a unimodal distribution. Therefore, we reparameterize that TPDM predicts two real numbers $a$ and $b$, and determines the distributions with Eq. 4. Thus the decay rate $r_{n}$ and the next noise level $t_{n}$ can be sampled as in Eq. 5-Eq. 6.
$\alpha=1+e^{a},\qquad\beta=1+e^{b},$ (4)
$r\sim\text{Beta}(\alpha,\beta)$ (5)
$t_{n}=r_{n}\cdot t_{n-1}$ (6)
TPDM requires only a minimal modification to the original diffusion model: adding a lightweight Time Prediction Module (TPM) as shown in Fig. 3. This module takes the latents before and after the transformer blocks as inputs, so that both the original noisy inputs and the predicted results at this step are taken into consideration. After several convolution layers, TPM pools the latents into a single feature vector and predicts $a$ and $b$ with two linear layers. We also use an adaptive normalization layer [41] in TPM so that the model is aware of the current time embedding.
During training, we freeze the original diffusion model and only update the newly introduced TPM. Thus, the model learns to predict the next diffusion time while preserving the original capacity for image generation.
Figure 4: From left to right, the images generated by TPDM-FLUX1.0-dev progress from simple to complex. Our Time Prediction Module adaptively adjusts the generation schedule to suit the complexity of each generation target.
Figure 5: Our TPDM-SD3-Medium, when compared to the SD3-Medium with the recommended and equivalent number of steps, demonstrates superior detail processing ability and generation accuracy.
Figure 6: The effect of $\gamma$ on the average number of inference steps.
3.3 Training Algorithms
To train TPM, we need to roll out at least two denoising steps: predicting the next diffusion time with the first step, and denoising with this time at the next step. A naive method provides a noised image as input to the first step and trains the model with the reconstruction loss calculated at the second step. The gradients would backpropagate through the predicted $t_{n}$ to update TPM in the first step.
However, we found that the trained model tends to complete the denoising process in very few steps during inference, leading to poor image quality. We hypothesize that by supervising the loss calculated after two steps, the model learns to generate a fully denoised image after two steps, and stop intermediately to minimize the loss function. However, what matters is the final image generated after the whole diffusion reverse process, which is ignored in this method.
Therefore, we optimize TPM to maximize the image quality generated after the whole denoising process to achieve precise time prediction. The quality is measured by an image reward model. Considering the computation graph of whole inference is too deep for gradients to backprop, we train the model with Proximal Policy Optimization (PPO) [48], whose loss function is formulated as
$\mathcal{L}(\theta)=-\left[\frac{\pi_{\theta}(y|s)}{\pi_{\text{old}}(y|s)}\hat{A}(s,y)-\lambda\text{KL}[\pi_{{\text{ref}}}(\cdot|s),\pi_{\theta}(\cdot|s)]\right]$ (7)
Under our problem definition, $s=(c,\epsilon)$ denotes initial states including the input prompt $c$ and gaussian noise $\epsilon\sim\mathcal{N}(0,1)$; $\pi_{\theta}$ denotes the policy network, e.g. our TPDM model. $\pi_{\text{old}}$ denotes the old policy for sampling trajectories; $\pi_{\text{ref}}$ denotes a reference policy for regularization; $y$ denotes the action our policy takes, i.e., the scheduled time; and $\hat{A}(s,y)$ denotes the advantage the action $y$ has given the state $s$.
We will specify the above elements in the PPO including the action and advantage we use in the following.
Treat the whole schedule as an action
Usually, when the model makes a sequence of predictions, PPO regards every single prediction as an action and optimizes with them as a batch. Recently, RLOO [1] claims that when the reward signal only appears at the end of the sequence and the environment dynamics is fully deterministic, we can regard the whole sequence as an action with no harm to the performance. Thus for simplicity, the entire generation, including all predicted time in the schedule, is considered as a single action to optimize.
Thus, when computing expectation in Eq. 7, we consider the whole trajectory as a training sample in optimization. However, TPDM only outputs the distribution of each single time prediction, which we denote by $\pi^{(i)}_{\theta}$. By the chain rule, the probability of the entire generation can be calculated as the product of each prediction as in Eq. 8,
$\pi_{\theta}(t_{1},...,t_{N}|s)=\prod_{i=1}^{N}\pi^{(i)}_{\theta}(t_{i}|s,t_{1},...,t_{i-1}),$ (8)
where $N$ denotes the total number of the generation steps. In our model, each factor of the probabilistic policy network $\pi^{(i)}_{\theta}$ is computed through the TPM that outputs a distribution over the possible diffusion time $t_{i}$ for the next step.
Image reward discounted by the number of steps
When generating samples for policy gradient training, we obtain trajectories by generating images from Gaussian noise with the current TPDM policy. Since there exists no ground truth for these generated images, we choose ImageReward [62] to assign a reward score based solely on the final image in the last time. Considering we do not want a trivially large number of too many steps during generation, we apply a discount factor $\gamma<1$ to discount reward to intermediate diffusion steps and calculate the average over the total number of denoising steps $N$. The final reward for this trajectory is shown in Eq. 9 below,
$R(s,y)=\frac{1}{N}\sum_{k=1}^{N}\gamma^{k-1}\text{IR}(y,s),$ (9)
where IR denotes the image reward model.
We refer readers to RLOO [1] on how to compute the advantage $\hat{A}(s,y)$ directly from a batch $R(s,y)$ without a value model for optimization. This reward function would encourage TPDM to generate images with high quality, and reduce the number of diffusion steps for model efficiency as well. By adjusting $\gamma$, one can control how many diffusion steps tend to be applied to generate high-quality images. The smaller the value of $\gamma$, the fewer steps would tend to be allowed. We will elaborate more in Sec. 4.2.
4 Experiments
4.1 Implementation details
Dataset
We collect prompts to train our model. These prompts are generated with Florence-2 [61] and Llava-Next [28] to generate captions for the Laion-Art [54] and COYO-700M [5] datasets, and utilize these prompts to constitute our training set. We will elaborate on this in Appendix.
Training Configurations
We use the AdamW optimizer with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.99, a constant learning rate of 1e-5, and a maximum gradient norm of 1.0. Our TPM module typically requires only 200 sampled trajectories, 4 PPO epochs, and a batch size of 256 to effectively update its parameters.
4.2 Main Results
Dynamic Schedule for Different Images
In Fig. 4, we present images generated with different prompts and their corresponding schedules predicted by TPDM. When prompting TPDM with shorter and simpler prompts, there exists fewer objects and details in the generated image, thus the diffusion time decreases faster and reaches 00 in relatively fewer steps. On the contrary, when longer and more complex prompts are provided, the model needs to generate more visual details. Therefore, the diffusion time decreases slower to generate delicate details. In this case TPDM requires more denoising steps in the generation process.
Models Inference Steps FID CLIP-T Aesthetic v2 Pick Score HPSv2.1
SD3-Medium [10] 28 25.00 0.322 5.433 22.12 29.12
SD3-Medium 15 24.72 0.321 5.426 22.30 28.52
TPDM-SD3-Medium 15.28 25.26 0.322 5.445 22.33 29.59
SD3.5-Large [2] 28 23.29 0.318 5.487 22.81 30.85
SD3.5-Large 15 23.35 0.323 5.475 22.62 30.02
TPDM-SD3.5-Large 15.22 24.48 0.322 5.525 22.81 30.64
FLUX.1-dev [23] 28 29.09 0.308 5.622 23.03 31.94
FLUX.1-dev 15 29.10 0.306 5.613 22.86 31.14
TPDM-FLUX.1-dev 13.57 28.98 0.314 5.685 22.94 30.77
Table 1: Evaluation of models across various benchmarks, best, second best
Adjusting $\gamma$ for Different Number of Steps
$\gamma$ in Eq. 9 controls how the image reward discounts with more generation steps, influencing how fast diffusion time decays during the denoising process, thereby affecting the average number of denoising steps of our model.
As shown in Fig. 6, as we decrease $\gamma$ from 0.97 to 0.85, TPDM tends to decrease the diffusion time more rapidly, leading to fewer sampling steps, from 15.0 to 7.5. Additionally, when compared to the baseline (yellow line), TPDM (purple line) consistently achieves a significantly higher aesthetic score with the same number of inference steps, achieving a good balance between model efficiency and generation performance.
Visual Comparison
Our method achieves superior ability in generating fine-grained details. Images generated by TPDM have a more realistic laptop keyboard than both images from SD3-Medium in Fig. 5C, and the results in Fig.5D also demonstrate that TPDM can generate images without errors. This may due to that TPDM can adjust the diffusion schedule according to the denoising quality.
4.3 Quantitative Results
We apply TPM on several state-of-the-art diffusion models, including Stable Diffusion 3 Medium, Stable Diffusion 3.5 Large, and Flux 1.0 dev, and demonstrate how TPM could enhance their performance. We primarily evaluate metrics from two kinds: The first is objective metrics including FID, alignment with the given prompts (CLIP-T), and human preference scores (Aesthetic Score v2, and HPSv2.1). The second is directly comparing images generated from different models via user study.
Quantitative Metrics
We compared TPDM and the above-mentioned models in Tab. 1. While maintaining competitive performance. All these models can generate images in about half of their recommended steps on average.
Moreover, the metrics representing human preference improve the most. For example, by generating with just 15.28 steps on average, TPDM-SD3-Medium gets a 29.59 HPS score, +1.07 higher than Stable Diffusion 3 with similar steps, also +0.47 higher than the original 28-step results. This may attributes to the reward model we utilized to calculate advantage in optimization, leading to aesthetically pleasing images aligned with human preferences. Furthermore, since other scheduling-tuning methods only conduct experiments on classical diffusion models such as Stable Diffusion 1.5 [45], we conduct an apple-to-apple comparison with these algorithms as well, which is demonstrated in Appendix Table 4.
User Study
To better reflect human attitudes towards these models, we conduct a user study by inviting volunteers to compare images generated from different models and select the one they prefer.
Specifically, for each prompt, we provide two images from SD3-Medium generated with 15 and 28 steps separately, and image generated from TPDM-SD3-Medium. We invited 15 volunteers to evaluate images generated from 50 prompts. The result is shown in Tab. 2, indicating that our model can generate images that better align human preferences.
Models Inference Steps Win Rate
SD3-Medium 28 26.58%
SD3-Medium 15 16.40%
TPDM-SD3-Medium 15.28 47.25%
Table 2: User study based on Stable Diffusion 3 architecture
4.4 Ablation on Module Architectures
In this section, we ablate on the choice of the inputs of TPM. As shown in Tab. 3, taking features from both the first and last layer into TPM performs better than only taking either of them.
TPM inputs Steps FID CLIP-T PickScore Aes v2
First 2 Layers 16.20 24.81 0.321 22.18 5.400
Last 2 Layers 19.30 23.19 0.322 22.16 5.356
First &\&& Last Layers 15.28 25.26 0.322 22.33 5.445
Table 3: Ablation Study of inputs hidden states
5 Conclusions and Limitations
In this paper, we introduce the Time Prediction Diffusion Model (TPDM), a text-to-image diffusion model that has flexible denoising schedulers for different individual prompts. By introducing a Time Prediction Module, we effectively train TPDM by reinforcement learning guided by a reward model. Based on the current leading diffusion model architecture proposed in Stable Diffusion 3 Medium, we train a strong MM-DiT-based TPDM that shows competitive quantitative performances on multiple text-to-image generation benchmarks.
Despite its promising performances, TPDM has some limitations. For instance, in this paper, we only design a relatively simple architecture for TPM. Whether and how to improve such a module to obtain better performances remains unexplored. Second, we froze the parameters of the original model and employed our training method to update the parameters of the diffusion model, leading to improved results. This approach warrants further exploration.
Supplementary Material
Appendix A Dataset Details
In Section 4, we provide a brief overview of our training dataset. Below, we present a more detailed description of the dataset.
A.1 Dataset Source
We curated prompts from image captions in two high-quality datasets, Laion-art [54]and COYO-11M [6]. The images in Laion-art are recaptioned with Florence-2 [61], while COYO-11M retained its original captions.
To ensure diversity during training, we design a filtering pipeline to select a diverse subset of these prompts.
A.2 Filtering Pipeline
Figure 7: The filtering pipeline used to construct the training set.
Figure 8: The statistics of prompts from the re-captioned high-resolution Laion-Art dataset, before and after filtering, highlight an improvement in diversity. Our dataset demonstrates greater variety compared to the original.
Our filtering pipeline, depicted in Figure 7, aims to select diverse and high-quality prompts by analyzing key linguistic features, such as nouns, prepositions, and adjectives. To assess diversity, we utilize WordNet [38] to count the number of valid nouns, adjectives, and prepositions in each prompt.
Prompts are ranked based on these counts. Those exhibiting the highest scores are included in the training set. The selection process prioritizes noun-diversity, ensuring a balanced representation across categories such as Person, Animal, Plant, Artifacts (Large/Small Objects), and Natural Views. Next, prompts with high preposition-diversity are selected, emphasizing those that contain spatial and relational terms (e.g., near, on). Finally, prompts are evaluated for adjective-diversity, with particular focus on adjectives describing color and shape, to enhance descriptive richness.
The process is iterative, selecting the top prompts in order of their noun, adjective, and preposition diversity scores until the desired data quantity (51,200 prompts, as specified in Section 4.1) and diversity are achieved. As shown in Figure 8, the diversity of the training set has significantly enhanced after filtering.
Appendix B Prompts of Figure
B.1 Figure 1
  1. 1.
    8k uhd A man looks up at the starry sky, lonely and ethereal, Minimalism, Chaotic composition Op Art.
  2. 2.
    A deep-sea exploration vessel descending into the pitch-black ocean, its powerful lights illuminating the glowing, alien creatures that inhabit the abyss. A massive, ancient sea creature with bioluminescent patterns drifts into view, its eyes glowing as it watches the explorers from the shadows of an underwater cave.
  3. 3.
    Half human, half robot, repaired human.
  4. 4.
    A baby painter trying to draw very simple picture, white background.
  5. 5.
    an astronaut sitting in a diner, eating fries, cinematic, analog film.
  6. 6.
    Van Gogh painting of a teacup on the desk.
  7. 7.
    A galaxy scene with stars, planets, and nebula clouds.
  8. 8.
    A hidden, forgotten city deep in a jungle, with crumbling stone temples overgrown by thick vines. In the heart of the city, a mysterious glowing artifact lies on an ancient pedestal, surrounded by an eerie mist. Strange symbols shimmer faintly on the stone walls, waiting to be uncovered.
  9. 9.
    A lone astronaut stranded on a desolate planet, gazing up at the sky. The planet’s surface is cracked and barren, with glowing, unearthly ruins scattered across the horizon. In the distance, a massive, alien ship slowly descends, casting an eerie shadow over the landscape.
Method Inference Steps FID CLIP-T Aesthetic v2 Pick Score HPSv2.1
DPM-Solver++ [30] 10 21.08 0.314 5.219 21.30 24.00
DPM-Solver++ & AYS [46] 10 18.81 0.313 5.156 21.31 24.40
DPM-Solver++ & GITS [7] 10 18.32 0.313 5.108 21.19 23.72
DPM-Solver++ & TPM 9.890 18.31 0.313 5.118 21.32 24.29
Table 4: Experiments on Stable Diffusion v1.5 [45]. 9.890 is the average step TPM used.
Appendix C Applicaion on Denoising Diffusion Probabilistic Models (DDPMs)
Prior to the widespread use of flow models, diffusion models dominates image generation, such as DDPM [13] and LDM [45]. They demonstrate several difference from flow models during inference. First, they directly predict the noise ($\epsilon$) added to the data at each step, rather than predicting the velocity vector field. Second, both the diffusion time $t$ and the noise level $\sigma_{t}$ are typically discrete, unlike continuous time in flow models. Moreover, the noise level $\sigma_{t}$ does not exhibit a simple linear relationship with the diffusion time $t$.
In this section, we integrate TPM into Stable Diffusion 1.5 [45], demonstrating that TPM also works in DDPMs. When determining the next diffusion time, TPM predicts the Beta distribution as in flow models, and then quantizes it to the closest $t$ in all $T$ discrete steps, thereby obtaining the corresponding noise level $\sigma_{t}$.
We can generally express their relationship as follows:
$\beta_{t}=\left(\sqrt{\beta_{start}}+\frac{t}{T-1}\left(\sqrt{\beta_{end}}-\sqrt{\beta_{start}}\right)\right)^{2},$ (10)
$\alpha_{t}=1-\beta_{t},$ (11)
$\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s},$ (12)
$\sigma_{t}=\sqrt{\frac{1-\bar{\alpha}_{t}}{\bar{\alpha}_{t}}},$ (13)
where $\beta_{t}$ is the variance of the noise added at diffusion time $t$. $\alpha_{t}$ is the proportion of the original signal remaining. $\sigma_{t}$ is the total noise level, and $T$ is the total number of diffusion steps while training, typically set to 1000. $\beta_{start}$ and $\beta_{end}$ are both hyper-parameter, controlling the initial and final noise variances.
We compare the performance of TPM with other scheduled fine-tuning methods applied to diffusion models. The evaluation metrics are calculated using 5,000 prompts from the COCO 2017 validation set, adhering to the evaluation protocol outlined in Table 1. Notably, our findings reveal that TPM can be effectively combined with higher-order solvers, such as DPM-Solver++ [30], resulting in a significant improvement of -2.79 in FID, surpassing the performance of other methods. It is important to note that GITS [7] requires access to the statistics of the evaluation set. While this approach is advantageous for large-scale image generation, it may offer limited benefits when generating a single image. Despite that GITS makes use of additional information, TPM consistently outperforms GITS across all evaluation metrics.